# Actively Learning Combinatorial Optimization Using a Membership Oracle

### v2

@@ 1 Introduction
In this paper, we take a step inside the development of specific techniques for combinatorial optimization in an interactive setting. We are interested in solving combinatorial optimization problems that are only partially described while learning a representation of the non-described portion. In order to limit the computational complexity of the optimization task, we assume a single unknown knapsack constraint. At the same time, we rely on a weak source of feedback consisting in receiving yes/no feasibility responses on single solutions. More formally, we want to solve:
@@
where weights $\bar{\boldsymbol{w}}\in W\subseteq[0,1]^{d}$ are unknown, the set $X$ encodes the known combinatorial structure of the problem in addition to the integrality constraints, while $W$ represents the prior knowledge at our disposal on the unknown weights $\bar{\boldsymbol{w}}$. We assume $W$ to be a polyhedron, i.e., $W=\{\boldsymbol{w}\in\mathbb{R}^{d}\ |\ \boldsymbol{a}\cdot\boldsymbol{w}\leq b\ \ \forall(\boldsymbol{a},b)\in\Delta\}\subseteq[0,1]^{d}$, where $|\Delta|<+\infty$. The unknown constraint (1b) is characterized by a membership oracle (with respect to $X$) that can determine the feasibility of any given solution $\boldsymbol{x}\in X$ with absolute certainty. Non-empty initial sets $S^{+}\subseteq X$ and $S^{-}\subseteq X$ are given, containing feasible and infeasible solutions to the unknown constraint (1b), respectively. We are also given a maximum number $N$ of oracle calls to respect, which reflects the fact that, when solving oracle-based problems in real contexts, the practical cost of a single call often plays a significant role.
@@
Inspired by active learning methods for binary classification based on Support Vector Machines (SVMs) [22], we define an iterative solution paradigm prescribing some key steps. The first one is a linear separation step that consists in training a linear classifier to separate the known feasible from the known infeasible solutions and to find the equation of the separating hyperplane $h:\hat{\boldsymbol{w}}\cdot\boldsymbol{x}=1$, with $\hat{\boldsymbol{w}}\in W$. In a second phase, we select an unlabeled solution applying a suitable sampling strategy and solve the 0-1 integer programming (IP) model using the surrogate knapsack constraint $\hat{\boldsymbol{w}}\cdot\boldsymbol{x}\leq 1$ in place of the unknown one. Both the solution selected by the sampling strategy and the optimal solution of the surrogate IP model are submitted to the oracle to be labeled. We emphasize that, from a theoretical standpoint, the only way to have a guarantee of optimality on a specific solution would be enumerating an exponential number of solutions, so, practically, the goal of our algorithmic framework is to compute the best solution, say $\hat{x}$, within the maximum number of oracle calls.
@@
We discuss an established approach in active learning that consists in using a SVM for linear classification together with a related sampling strategy that is function of the current separating hyperplane (a so-called information-based strategy). In particular, we employ a strategy known as simple margin [37], which applies to our combinatorial setting in a straightforward way. We propose an alternative sampling strategy called closest cutting plane that exploits the discrete nature of the problem and the structure of the so-called version space, i.e., the set of all surrogate weights $\hat{\boldsymbol{w}}\in W$ such that the hyperplane $h:\hat{\boldsymbol{w}}\cdot\boldsymbol{x}=1$ linearly separates the current labeled solutions. Subsequently, we propose an alternative linear separation method obtained as a variant of an algorithm for convex optimization with a separation oracle [8]. [^1] By separation in the separation oracle, we mean finding a hyperplane that separates a point from a convex set, which is different from separation in the linear separation of labeled points. Given a convex set $K\subseteq\mathbb{R}^{d}$, a separation oracle for the set takes in input any $\boldsymbol{w}\in\mathbb{R}^{d}$ and outputs either “yes” if $\boldsymbol{w}\in K$ or a linear inequality valid for $K$ but violated by $\boldsymbol{w}$. Our version of the algorithm computes the vector of surrogate weights relying exclusively on the membership oracle at our disposal, which is used to approximate a separation oracle.
We conduct experiments to assess the effectiveness of different combinations of linear separation method and sampling strategy to approximate an optimal binary solution. For this purpose, we consider two specific realizations of the original problem (1). The first one is the pure knapsack case, in which $X=\{0,1\}^{d}$. The second one is inspired by the generation of personalized learning systems with uncertain item weights [1] and features known constraints in addition to the unknown knapsack constraint.
@@
The remainder of the paper is organized as follows. In Section 2, we give some background and we review the related literature. In Section 3, we give the details of our solution approaches. In Section 4, we present our experimental setting and some empirical results. Finally, we draw our conclusions in Section 5.
@@ 2 Background and Literature Review
In this Section we review the literature and give some background about convex optimization in the oracle setting, then we briefly present some principles and methodologies of active learning on which our algorithmic contributions rely heavily.
@@ 2.1 Oracle-Based Optimization
Oracle-based optimization has received substantial attention over the years for its role in applications where an explicit description of the feasible region is not available or difficult to formalize. A well-studied setting consists in minimizing a convex function $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$ over a convex set $K$ equipped with a separation oracle [3, 7, 8, 13, 14, 21, 25, 29, 39], which is proven to be solvable with a pseudo-linear number of oracle calls with respect to the dimension $d$. Recent advances include convex optimization with a membership oracle of $K$ and an evaluation oracle of $f$ [24]. In the case of convex sets, separation and membership oracles are equivalent from a complexity standpoint. In fact, according to the classical result in [15], there exists a polynomial time equivalence between optimization, separation and membership for convex bodies. Since in this paper we are interested in the combinatorial optimization scenario, convenient theoretical results like those described above are not available due to the non-convexity of the feasible region. Actually, as shown in the example of Appendix A, any randomized algorithm requires $\Omega(n^{1/\epsilon})$ calls to the membership oracle in expectation to find a $(1-\epsilon)$ optimal solution of a knapsack problem with an unknown knapsack constraint in the worst case.
@@ 2.2 Active Learning for Linear Separation
The solution approaches that we propose in this paper to solve the general problem introduced in Section 1 are based on learning a surrogate linear constraint or, in other terms, an approximation of the halfspace containing the feasible points according to the original unknown constraint. Halfspace learning is a research area with a long tradition that traces back to the Perceptron algorithm [32]. It has received extensive interest over the years (see, e.g., [2, 4, 5, 9, 10, 17, 35]), especially regarding popular algorithms like, e.g., Support Vector Machines (SVMs) [6, 27, 40]. A significant part of the work in halfspace learning has been done in the context of active learning [12, 18, 23, 34, 36]. Compared to traditional supervised learning techniques, that can usually require massive amounts of labeled data to produce models with the desired generalization capabilities, active learning algorithms employ relatively small initial labeled datasets and are based on querying an oracle on new unlabeled data in order to make the learning process less data consuming [36]. A particular sub-category of active learning methods is represented by algorithms whose distinctive characteristic is to rely on a fixed dataset of unlabeled data and to implement a policy, called sampling strategy, to decide what unlabeled available points submit to the oracle for labeling, in a sequential fashion. An established active learning method for binary classification consists in alternating the training of a binary classifier on the dataset of already labeled points and a sampling step. The classification task is traditionally accomplished using a SVM [19, 33, 38, 41], while several related sampling strategies have been proposed in the literature (see, e.g., [22] and [36]). Existing strategies fall in one of three categories: information-based strategies, that sample the most informative points, for example selecting those whose label is the most uncertain; representation-based strategies, that try to diversify the areas where the points are sampled; and finally, strategies that combine the two previous approaches.
In the presence of separable data, the set of all the classifiers in the hypothesis space that correctly separate the current observations is called version space [28, 36]. Training a classifier on the current labeled points has the “dual” meaning of finding a point inside the current version space. Sampling an unlabeled point and acquiring its label from the oracle corresponds to constraint and possibly reduce the version space in size. The goal behind the idea of alternating the training of a binary classifier and a sampling strategy is to find a convenient “dual” point inside the version space and subsequently to sample a point in the original “primal” space whose effect is to significantly shrink the version space.
@@ 3 Active Learn-and-Optimize
Algorithm 1 Active learn-and-optimize
2:$t\leftarrow 0$
3:$n_{q}\leftarrow 0$
4:while $n_{q}<N$ do
5:     $\hat{\boldsymbol{w}}_{t}\leftarrow\mathcal{S}(W,S^{+}_{t},S^{-}_{t})$triangleright Linear separation
6:     $t\leftarrow t+1$
7:     $P^{+}_{t}\leftarrow\emptyset$
8:     $P^{-}_{t}\leftarrow\emptyset$
9:     for $\theta\in\{\text{sampling},\text{optimization}\}$ do
10:         if $\theta=\text{sampling}$ then
11:              $\boldsymbol{u}_{t}\leftarrow\mathcal{T}(X,S^{+}_{t-1},S^{-}_{t-1})$triangleright Sampling strategy
12:         else
13:              Get $\boldsymbol{u}_{t}\in\arg\max\left\{\boldsymbol{v}\cdot\boldsymbol{x}\ |\ \boldsymbol{x}\in X\setminus(S^{+}_{t-1}\cup S^{-}_{t-1})\text{,}\ \hat{\boldsymbol{w}}_{t}\cdot\boldsymbol{x}\leq 1\right\}$triangleright Optimization
14:         $\ell_{\boldsymbol{u}_{t}}\leftarrow\Omega(\boldsymbol{u}_{t})$triangleright Oracle call
15:         $n_{q}\leftarrow n_{q}+1$
16:         if $\ell_{\boldsymbol{u}_{t}}=1$ then
17:              $P^{+}_{t}\leftarrow P^{+}_{t}\cup\{\boldsymbol{u}_{t}\}$
18:         else
19:              $P^{-}_{t}\leftarrow P^{-}_{t}\cup\{\boldsymbol{u}_{t}\}$
20:         if $n_{q}=N$ then $\triangleright$ Stopping criterion
21:              break
22:     $S^{+}_{t}\leftarrow S^{+}_{t-1}\cup P^{+}_{t}$
23:     $S^{-}_{t}\leftarrow S^{-}_{t-1}\cup P^{-}_{t}$
24:Get $\hat{\boldsymbol{x}}\in\arg\max\left\{\boldsymbol{v}\cdot\boldsymbol{x}\ |\ \boldsymbol{x}\in X,\boldsymbol{x}\leq\boldsymbol{y}\text{ for some }\boldsymbol{y}\in S^{+}_{t}\right\}$triangleright Final solution
25:$\hat{\boldsymbol{w}}\leftarrow\mathcal{S}(W,S^{+}_{t},S^{-}_{t})$triangleright Final linear separation
26:return $\hat{\boldsymbol{x}}$, $\hat{\boldsymbol{w}}$
We can conceptually identify a few main steps inside the framework: the linear separation step (line 5) finds the surrogates weights by computing $\mathcal{S}(W,S^{+}_{t},S^{-}_{t})$; the sampling step (line 11) is the application of the chosen sampling strategy, i.e., $\mathcal{T}(X,S^{+}_{t},S^{-}_{t})$, to select an unlabeled point; the optimization step (line 13) finds an optimal solution of the partially specified 0-1 IP model using the current surrogate constraint in place of the unknown knapsack one. Both the sampling and the optimization steps are followed by an oracle call that returns the label of the newly found point (line 14). The stopping criterion (line 20) is triggered by reaching the maximum number of queries. In cases when $\boldsymbol{v}\geq 0$, one can simplify the final solution step (line 24) to $\hat{\boldsymbol{x}}\in\arg\max\left\{\boldsymbol{v}\cdot\boldsymbol{x}\ |\ \boldsymbol{x}\in S^{+}_{t}\right\}$. We remark that, since the sampling step and the optimization step always return new points with respect to the previous iterations, and since the set $X$ is finite, any instance of the framework ensures finite convergence to the optimal solution when permitting an infinite number of oracle calls. Finally, we observe that, in principle, the framework admits modifications such as employing multiple sampling or optimization steps between any two consecutive linear separation steps, or selecting just one single point to query at every time step $t$.
As is common in active learning, SVM can be used to accomplish the linear separation task. Specifically, we may solve the following linear programming model and set $\boldsymbol{w}/b$ as the surrogate weights:
@@
$\ \boldsymbol{w}\in\mathbb{R}^{d},~{}b\in\mathbb{R}.$
For the sampling step, one can consider multiple existing approaches. However, established SVM-related strategies like, e.g., MaxMin margin or Ratio margin, require to train a linear classifier for every unlabeled point [22]. In general, most of the sampling strategies present in the literature are incompatible with a combinatorial optimization setting, in which there is an exponential number of unlabeled examples to choose from. To the best of our knowledge, the only existing strategy that can be efficiently adapted to such a setting is Simple margin (SIM) [37]. It consists in selecting the point that maximizes the uncertainty with respect to the classification provided by the current SVM, i.e., the unlabeled point closest to the decision boundary. The task can be accomplished by solving the following model:
@@
The objective function (2a) minimizes the distance between the generic point $\boldsymbol{u}\in X$ and the hyperplane identified by $\hat{\boldsymbol{w}}$. No-good cuts (2b) and (2c) are used to prevent finding a solution whose feasibility or infeasibility is implied by labeled points in the sets $S^{+}_{t}$ and $S^{-}_{t}$. We remark that, although integrality constraints make the problem NP-hard in the general case, mixed-integer optimization models can be often solved efficiently, e.g., by off-the-shelf discrete optimization solvers.

### v3

@@ Abstract
Keywords:
Combinatorial optimization Interactive optimization Active learning.
@@ 1 Introduction
In this paper, we take a step inside the development of specific techniques for combinatorial optimization in an interactive setting. We are interested in solving combinatorial optimization problems that are only partially described while learning a representation of the non-described portion. In order to limit the computational complexity of the optimization task, we assume a single unknown knapsack constraint. At the same time, we rely on a weak source of feedback consisting in receiving yes/no feasibility responses on single solutions. More formally, we want to solve
@@
where the set $X$ encodes the known combinatorial structure of the problem in addition to the integrality constraints, weights $\bar{\boldsymbol{w}}\in W\subseteq[0,1]^{d}$ are unknown, while $W$ represents the prior knowledge at our disposal on the unknown weights $\bar{\boldsymbol{w}}$. We assume $W$ to be a polyhedron, i.e., $W=\{\boldsymbol{w}\in\mathbb{R}^{d}\ |\ \boldsymbol{a}\cdot\boldsymbol{w}\leq b\ \ \forall(\boldsymbol{a},b)\in\Delta\}\subseteq[0,1]^{d}$, where $|\Delta|<+\infty$. The unknown constraint (1b) is characterized by a membership oracle (with respect to $X$) that can determine the feasibility of any given solution $\boldsymbol{x}\in X$ with absolute certainty. We assume initial sets $S^{+}\subseteq X$ and $S^{-}\subseteq X$ are given, containing feasible and infeasible solutions to the unknown constraint (1b), respectively. We are also given a maximum number $N$ of oracle calls to respect, which reflects the fact that, when solving oracle-based problems in real contexts, the practical cost of a single call often plays a significant role.
From a theoretical standpoint, the oracle complexity of optimizing problem (1) is exponential. The proof of the following result is in Appendix 0.A.
Theorem 1.1
For any deterministic algorithm for Problem (1), for any $\epsilon\in(0,1)$, there exists an instance of the problem that takes $\Omega(d^{1/\epsilon})$ oracle calls to label a feasible solution whose objective value is within a multiplicative factor of $(1-\epsilon)$ with respect to the optimal solution.
@@
Given such hardness result, our goal is to find a practical algorithm that computes the best solution to the problem, say $\hat{x}$, within the given number of oracle calls. Inspired by active learning methods for binary classification based on Support Vector Machines (SVMs) [22], we define an iterative solution paradigm prescribing some key steps. The first one is a linear separation step that consists in training a linear classifier to separate the known feasible from the known infeasible solutions and to find the equation of the separating hyperplane $h:\hat{\boldsymbol{w}}\cdot\boldsymbol{x}=1$, with $\hat{\boldsymbol{w}}\in W$. In a second phase, we apply a suitable sampling strategy and solve the 0-1 integer programming (IP) model using the surrogate knapsack constraint $\hat{\boldsymbol{w}}\cdot\boldsymbol{x}\leq 1$ in place of the unknown one to select solutions. Both the solution selected by the sampling strategy and the optimal solution of the surrogate IP model are submitted to the oracle to be labeled.
@@
We discuss an established approach in active learning that consists in using a SVM for linear classification together with a related sampling strategy that is function of the current separating hyperplane (a so-called information-based strategy). In particular, we employ a strategy known as simple margin [37], which applies to our combinatorial setting in a straightforward way. We propose an alternative sampling strategy called closest cutting plane that exploits the discrete nature of the problem and the structure of the so-called version space, i.e., the set of all surrogate weights $\hat{\boldsymbol{w}}\in W$ such that the hyperplane $h:\hat{\boldsymbol{w}}\cdot\boldsymbol{x}=1$ linearly separates the current labeled solutions. Subsequently, we propose an alternative linear separation method obtained as a variant of an algorithm for convex optimization with a separation oracle [8]. [^1] By separation for a separation oracle, we mean finding a hyperplane that separates a point from a convex set, which differs from the notion of separation in linear separation of labeled points. Given a convex set $K\subseteq\mathbb{R}^{d}$, a separation oracle for the set takes in input any $\boldsymbol{w}\in\mathbb{R}^{d}$ and outputs either “yes” if $\boldsymbol{w}\in K$ or a linear inequality valid for $K$ but violated by $\boldsymbol{w}$. Our version of the algorithm computes the vector of surrogate weights relying exclusively on the membership oracle at our disposal, which is used to approximate a separation oracle.
@@
The remainder of the paper is organized as follows. In Section 2, we give some background and we review the related literature. In Section 3, we give the details of our solution approaches. In Section 4, we present our experimental setting and some empirical results on a knapsack problem and a college study plan problem. Finally, we draw our conclusions in Section 5.
@@ 2.1 Oracle-Based Optimization
Oracle-based optimization has received substantial attention over the years for its role in applications where an explicit description of the feasible region is not available or difficult to formalize. A well-studied setting consists in minimizing a convex function $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$ over a convex set $K$ equipped with a separation oracle [3, 7, 8, 13, 14, 21, 25, 29, 39], which is proven to be solvable with a pseudo-linear number of oracle calls with respect to the dimension $d$. Recent advances include convex optimization with a membership oracle of $K$ and an evaluation oracle of $f$ [24]. In the case of convex sets, separation and membership oracles are equivalent from a complexity standpoint. In fact, according to the classical result in [15], there exists a polynomial time equivalence between optimization, separation and membership for convex bodies. Since in this paper we are interested in the combinatorial optimization scenario, convenient theoretical results like those described above are not available due to the non-convexity of the feasible region. Actually, as shown in Theorem 1.1, any algorithm requires $\Omega(d^{1/\epsilon})$ calls to the membership oracle in expectation to find a $(1-\epsilon)$-approximate solution of a knapsack problem with an unknown knapsack constraint in the worst case.
@@ 2.2 Active Learning for Linear Separation
The solution approaches that we propose in this paper to solve the general problem introduced in Section 1 are based on learning a surrogate linear constraint or, in other terms, an approximation of the halfspace containing the feasible points according to the original unknown constraint. Halfspace learning is a research area with a long tradition that traces back to the Perceptron algorithm [32]. It has received extensive interest over the years (see, e.g., [2, 4, 5, 9, 10, 17, 35]), especially regarding popular algorithms like, e.g., Support Vector Machines (SVMs) [6, 27, 40]. A significant part of the work in halfspace learning has been done in the context of active learning [12, 18, 23, 34, 36]. Compared to traditional supervised learning techniques that can usually require massive amounts of labeled data to produce models with the desired generalization capabilities, active learning algorithms employ relatively small initial labeled datasets and are based on querying an oracle on new unlabeled data in order to make the learning process less data consuming [36]. A particular sub-category of active learning methods is represented by algorithms whose distinctive characteristic is to rely on a fixed dataset of unlabeled data and to implement a policy, called sampling strategy, to decide what unlabeled available points submit to the oracle for labeling, in a sequential fashion. An established active learning method for binary classification consists in alternating the training of a binary classifier on the dataset of already labeled points and a sampling step. The classification task is traditionally accomplished using a SVM [19, 33, 38, 41], while several related sampling strategies have been proposed in the literature (see, e.g., [22] and [36]). Existing strategies fall in one of three categories: information-based strategies, that sample the most informative points, for example selecting those whose label is the most uncertain; representation-based strategies, that try to diversify the areas where the points are sampled; and finally, strategies that combine the two previous approaches.
In the presence of separable data, the set of all linear classifiers that correctly separate the current observations is called version space [28, 36]. Training a classifier on the current labeled points means finding a point inside the current version space. Sampling an unlabeled point and acquiring its label from the oracle corresponds to possibly reducing the version space in size, since every new labeled point either restricts the set of classifiers that can separate the current labeled data or leaves it unchanged. The idea behind alternating between training a binary classifier and sampling is that a convenient combination of the two steps can shrink the version space significantly, thus facilitating the convergence towards better classifiers for the binary classification problem.
@@ 3 Active Learn-and-Optimize
We can conceptually identify a few main steps inside the framework: the linear separation step (line 5) finds the surrogates weights by computing $\mathcal{S}(W,S^{+}_{t},S^{-}_{t})$; the sampling step (line 9) is the application of the chosen sampling strategy, i.e., $\mathcal{T}(X,S^{+}_{t},S^{-}_{t})$, to select an unlabeled point; the optimization step (line 11) finds an optimal solution of the partially specified 0-1 IP model using the current surrogate constraint in place of the unknown knapsack one. Both the sampling and the optimization steps are followed by an oracle call that returns the label of the newly found point (line 13). The stopping criterion (line 18) is triggered by reaching the maximum number of queries. In cases when $\boldsymbol{v}\geq 0$, one can simplify the final solution step (line 21) to $\hat{\boldsymbol{x}}\in\arg\max\left\{\boldsymbol{v}\cdot\boldsymbol{x}\ |\ \boldsymbol{x}\in S^{+}_{t}\right\}$. We remark that, since the sampling step and the optimization step always return new points with respect to the previous iterations, and since the set $X$ is finite, any instance of the framework ensures finite convergence to the optimal solution when permitting an infinite number of oracle calls. Finally, we observe that, in principle, the framework admits modifications such as employing multiple sampling or optimization steps between any two consecutive linear separation steps, or selecting just a single point to query at every time step $t$.
As is common in active learning, SVM can be used to accomplish the linear separation task. Specifically, we may solve the following quadratic programming model and set $\boldsymbol{w}/b$ as the surrogate weights:
@@
$\ \boldsymbol{w}\in W,~{}b\in\mathbb{R}.$
Algorithm 1 Active learn-and-optimize
2:$t\leftarrow 0$, $n_{q}\leftarrow 0$
3:while $n_{q}<N$ do
4:     $t\leftarrow t+1$
5:     $\hat{\boldsymbol{w}}_{t}\leftarrow\mathcal{S}(W,S^{+}_{t-1},S^{-}_{t-1})$triangleright Linear separation
6:     $P^{+}_{t}\leftarrow\emptyset$, $P^{-}_{t}\leftarrow\emptyset$
7:     for $\theta\in\{\text{sampling},\text{optimization}\}$ do
8:         if $\theta=\text{sampling}$ then
9:              $\boldsymbol{u}_{t}\leftarrow\mathcal{T}(X,S^{+}_{t-1},S^{-}_{t-1})$triangleright Sampling strategy
10:         else
11:              Get $\boldsymbol{u}_{t}\in\arg\max\left\{\boldsymbol{v}\cdot\boldsymbol{x}\ |\ \boldsymbol{x}\in X\setminus(S^{+}_{t-1}\cup S^{-}_{t-1})\text{,}\ \hat{\boldsymbol{w}}_{t}\cdot\boldsymbol{x}\leq 1\right\}$
12:              $\triangleright$ Optimization
13:         $\ell_{\boldsymbol{u}_{t}}\leftarrow\Omega(\boldsymbol{u}_{t})$, $n_{q}\leftarrow n_{q}+1$triangleright Oracle call
14:         if $\ell_{\boldsymbol{u}_{t}}=1$ then
15:              $P^{+}_{t}\leftarrow P^{+}_{t}\cup\{\boldsymbol{u}_{t}\}$
16:         else
17:              $P^{-}_{t}\leftarrow P^{-}_{t}\cup\{\boldsymbol{u}_{t}\}$
18:         if $n_{q}=N$ then $\triangleright$ Stopping criterion
19:              break
20:     $S^{+}_{t}\leftarrow S^{+}_{t-1}\cup P^{+}_{t}$, $S^{-}_{t}\leftarrow S^{-}_{t-1}\cup P^{-}_{t}$
21:Get $\hat{\boldsymbol{x}}\in\arg\max\left\{\boldsymbol{v}\cdot\boldsymbol{x}\ |\ \boldsymbol{x}\in X,\boldsymbol{x}\leq\boldsymbol{y}\text{ for some }\boldsymbol{y}\in S^{+}_{t}\right\}$triangleright Final solution
22:$\hat{\boldsymbol{w}}\leftarrow\mathcal{S}(W,S^{+}_{t},S^{-}_{t})$triangleright Final linear separation
23:return $\hat{\boldsymbol{x}}$, $\hat{\boldsymbol{w}}$
For the sampling step, one can consider multiple existing approaches. However, established SVM-related strategies like, e.g., max-min margin or ratio margin, require to train a linear classifier for every unlabeled point [22]. In general, most of the sampling strategies in the literature are incompatible with a combinatorial optimization setting, in which there is an exponential number of unlabeled examples to choose from. To the best of our knowledge, the only existing strategy that can be efficiently adapted to such a setting is simple margin (SIM) [37]. At every iteration $t$, it consists in selecting the point that maximizes the uncertainty with respect to the classification provided by the current SVM, i.e., the unlabeled point closest to the decision boundary. The task can be accomplished by solving the following optimization problem:
@@
The objective function (2a) minimizes the distance between the generic point $\boldsymbol{u}\in X$ and the hyperplane identified by $\hat{\boldsymbol{w}}$. No-good cuts (2b) and (2c) are used to prevent finding a solution whose feasibility or infeasibility is implied by labeled points in the sets $S^{+}_{t}$ and $S^{-}_{t}$. We remark that, although integrality requirements (2d) make the problem NP-hard in the general case, mixed-integer optimization models can be often solved efficiently, e.g., by off-the-shelf discrete optimization solvers.